Two forces parallel to the x axis do 19.0 J of work on a small tray while moving it 17.6 m in the x direction across a gym floor. One of the forces has a value of +3.80 N in the x direction. What is the other force?
Answer #1

Work = Force * Distance


Work = (F2-F1) * Distance


19 = (F2+3.8) * 17.6


F2 = - 2.72 N(opp direction)
Answer #2

19=(3.80 +F)*17.6


F=-2.72 N